# PTDiffusion: Free Lunch for Generating Optical Illusion Hidden Pictures with Phase-Transferred Diffusion Model

### v2

@@ Abstract
Optical illusion hidden picture is an interesting visual perceptual phenomenon where an image is cleverly integrated into another picture in a way that is not immediately obvious to the viewer. Established on the off-the-shelf text-to-image (T2I) diffusion model, we propose a novel training-free text-guided image-to-image (I2I) translation framework dubbed as Phase-Transferred Diffusion Model (PTDiffusion) for hidden art syntheses. PTDiffusion embeds an input reference image into arbitrary scenes as described by the text prompts, while exhibiting hidden visual cues of the reference image. At the heart of our method is a plug-and-play phase transfer mechanism that dynamically and progressively transplants diffusion features’ phase spectrum from the denoising process to reconstruct the reference image into the one to sample the generated illusion image, realizing harmonious fusion of the reference structural information and the textual semantic information. Furthermore, we propose asynchronous phase transfer to enable flexible control to the degree of hidden content discernability. Our method bypasses any model training and fine-tuning, all while substantially outperforming related methods in image quality, text fidelity, visual discernibility, and contextual naturalness for illusion picture synthesis, as demonstrated by extensive qualitative and quantitative experiments.
@@ 1 Introduction
Since diffusion models [21] revolutionizing the field of generative AI, tremendous attention has been focused on various diffusion-based AIGC applications, among which there are also explorations in illusion picture synthesis. For example, DiffQRCoder [29] and Text2QR [47] leverage ControlNet [48] to integrate scannable QR codes into aesthetic pictures. Diffusion Illusions [2] employs T2I diffusion model and score distillation sampling [38, 46] to synthesize images with overlay illusions. Visual Anagrams [16] combines noises estimated from different views to generate multi-view optical illusions, realizing image appearance change under a certain pixel permutation such as image flip, image rotation, or jigsaw rearrangement.
@@
In this paper, we pioneer generating optical illusion hidden pictures (we will use “illusion pictures” as an abbreviation in the following) from the perspective of text-guided I2I translation, i.e., translating an input reference image into an illusion picture that complies with the text prompt in semantic content while manifesting structural visual cues of the reference image. Our goal differs from the aforementioned optical illusion methods in three aspects: (i) different from camouflage image generation [3, 49, 28] that overemphasizes content concealment, we pursue visual discernibility of both target semantic content and original hidden content; (ii) unlike synthesizing camouflage image that conceals content into the texture of an existing background image, we expect generating background as per the text; (iii) we do not aim at producing transformation-based (flip, rotation, etc.) optical illusions like Diffusion Illusions [2] and Visual Anagrams [16], but rather focus on seamlessly dissolving a reference image into arbitrary scenes. By contrast, our goal is more methodologically relevant to text-guided I2I [31, 45, 36] and controllable T2I [48, 32, 50] methods. However, since these methods over-bind I2I correlation by explicitly enforcing feature consistency constraints [31, 45, 36] or directly training a control network [48, 32], they are less suitable for I2I translation with large semantic deviation (i.e., suffer from structure-semantic conflict issue), and thus tend to generate contextually unnatural results when applied to synthesize illusion pictures. This enlightens us to explore a disentangled image structure representation to relax I2I correlation binding, as well as an appropriate manner to deeply fuse image structure and semantic information along the T2I sampling process.
@@
Drawing inspiration from digital signal processing that the phase spectrum of an image determines its structural composition, we initiatively propose to leverage diffusion features’ phase to disentangle image structure and accordingly propose PTDiffusion, a concise and elegant method based on T2I diffusion model that realizes smooth blending of the reference image’s structural cues and the text-instructed semantic content in the Latent Diffusion Model (LDM) [40] feature space, producing visually appealing illusion pictures in a plug-and-play manner. Specifically, we employ DDIM inversion [43] to extract guidance features along a reference image reconstruction trajectory, and progressively transplant the phase of the guidance features into the corresponding features along the text-guided sampling trajectory, such that structural cues of the input reference image are smoothly penetrated into the sampling process of the target image, yielding generation results exhibiting harmonious illusion effects. Besides, we further propose asynchronous phase transfer to flexibly control the structural penetration strength, endowing our method with controllability to the hidden content discernability in the generated illusion picture. Our method is free from model training, fine-tuning, and online optimization, all while demonstrating noticeable strengths in illusion picture synthesis. The contributions of this work are summarized as follows:
@@ 2 Related work
Text-guided image generation. Since the advent of DDPM [21], diffusion model has soon surpassed GAN [17] on image synthesis [4] and has subsequently been accelerated by DDIM [43] and extended to conditional image generation paradigm by Palette [41]. After large-scale T2I diffusion models [33, 42, 39] remarkably boosting AIGC industry, LDM [40] contributes a classical T2I framework with dramatically lowered computational overhead by transferring DDPM from high-dimensional pixel space into low-dimensional feature space, which inspires subsequent T2I models [37, 7] scaling up to larger capacity. To add controllability to T2I synthesis, ControlNet [48] and T2I-Adapter [32] train a control network of the LDM conditioned on certain image priors (edges, depth maps, etc.), implicitly realizing reference-image-based structural control to the generated images. However, the over-constraining on object contours and shapes of these methods limit their applicability to synthesizing illusion pictures which emphasizes harmonious blending of image structure and semantics.
Image-to-image translation. I2I translation is initially handled by GAN [17] in a supervised manner [22], and is subsequently extended to unsupervised paradigm through cycle-consistency loss [51], geometry-consistency constraint [9], contrastive learning [35], etc., dramatically enriching image stylization applications such as painting stylization [12, 27, 11], font stylization [1, 23], and image cartoonization [13, 24]. Since the stunning success of T2I synthesis [33, 42, 39, 40], attention has been focused on extending large-scale T2I diffusion models to the realm of text-guided I2I translation. SDEdit [30] translates a reference image by noising it to an intermediate step followed by text-guided denoising. Attention-modulation-based methods such as Null-text inversion [31] and pix2pix-zero [36] establish I2I correlation by enforcing consistency of cross-attention maps [18]. Textual-inversion-based methods like Imagic [26] and Prompt Tuning Inversion [5] preserve source image visual information via learnable text embedding. Optimization-free methods represented by PAP [45] maintain source image structure through dynamic feature injection during the reverse sampling process. Besides, FCDiffusion [14] and FBSDiff [10] leverage DCT frequency bands of diffusion features to control different I2I correlation factors. For illusion picture synthesis, however, these methods struggle to produce contextually natural results that are both faithful to the target textual semantics and discernable of the source image structural cues.
@@ 3.1 Overall architecture
To naturally embed $x$ into $\tilde{x}$, we propose phase transfer module (PTM) which dynamically blends the structural information of $\hat{z}_{t}$ into $\tilde{z}_{t}$ along the two parallel denoising trajectories. We apply per-step structural penetration realized by PTM only in the early part of the sampling trajectory (which we call phase transfer stage) while leaving the rear part (which we call refining stage) totally unconstrained to guarantee high-quality image synthesis. The two stages are separated by the time step $\lambda T$, where $\lambda$ denotes the proportion of the refining stage to the entire sampling trajectory. The final sampling result $\tilde{z}_{0}$ is transformed to the generated illusion picture via the LDM decoder $D$, i.e., $\tilde{x}=D(\tilde{z}_{0})$.
@@ 3.2 Phase transfer module
where $i$ denotes the imaginary unit, i.e., $i^{2}=-1$. $\hat{R}_{t}$ and $\tilde{R}_{t}$ are the real part of the 2D FFT result of $\hat{z}_{t}$ and $\tilde{z}_{t}$ respectively, $\hat{I}_{t}$ and $\tilde{I}_{t}$ are the corresponding imaginary part. $\hat{M}_{t}$ and $\tilde{M}_{t}$ are the magnitude spectrum of $\hat{z}_{t}$ and $\tilde{z}_{t}$ respectively, while $\hat{P}_{t}$ and $\tilde{P}_{t}$ are the phase spectrum of them. Then, the calculated phase spectra $\hat{P}_{t}$ and $\tilde{P}_{t}$ are linearly blended with a time-dependent blending coefficient $b_{t}$ to yield the fused phase spectrum $P_{t}^{fuse}$:
@@
The fused phase is recombined with the original magnitude $\tilde{M}_{t}$ before transforming back to the spatial domain with 2D IFFT, finally yielding the structurally penetrated $\tilde{z}_{t}$:
@@
Since the structural information of the guidance features $\hat{z}_{t}$} is becoming increasingly prominent as the denoising process proceeds, direct phase transfer in the later denoising steps is prone to harm contextual naturalness of the final result due to excessive structural penetration. Thus, we design a decayed phase blending schedule that gradually decays the blending coefficient {$b_{t}$} in the later denoising steps of the phase transfer stage, which we formulate as below:
@@
where $\lambda\leq\tau\leq 1$. It means that we apply direct phase replacement at the early part of the phase transfer stage where the guidance features $\hat{z}_{t}$ are structurally less distinct, while gradually decaying phase transfer intensity in the later section of the phase transfer stage as $\hat{z}_{t}$ becomes structurally more and more prominent. By decaying phase transfer intensity to avoid overly penetrated structural information, the visual quality and harmony degree of the generated illusion picture can be noticeably improved.
@@ 3.3 Asynchronous phase transfer
Since sufficient denoising steps in the refining stage are of crucial importance to ensure high-quality image generation, we fix $\lambda$, namely the length of the phase transfer stage, and propose asynchronous phase transfer to realize controllable structural penetration strength within the fixed-length phase transfer stage. As Fig. 4 displays, later $\hat{z}_{t}$ in the reconstruction trajectory is structurally more prominent than its earlier counterpart. This inspires us to transfer phase from later $\hat{z}_{t}$ with smaller time steps into earlier $\tilde{z}_{t}$ with larger time steps to enhance structural penetration, namely the so-called asynchronous phase transfer as illustrated in Fig. 5.
@@
To this end, we design a concise and elegant solution that implements asynchronous phase transfer based on synchronous denoising, i.e., parallel denoising along the reconstruction and sampling trajectory. Specifically, given an async distance $d$, we firstly leverage $\hat{z}_{t}$ to estimate its future counterpart $\hat{z}_{t-d}$ that is $d$ steps ahead of $\hat{z}_{t}$ in the reconstruction trajectory, then transfer the phase spectrum of the estimated $\hat{z}_{t-d}$ into $\tilde{z}_{t}$. Let APTM denote the asynchronous phase transfer module, we formulate it as follows:
@@ 4.1 Qualitative results
As demonstrated in Fig. 7, despite using deterministic DDIM inversion and sampling, our method enables to sample diversified results simply by varying the initialized Gaussian noise $\tilde{z}_{T}$, while existing advanced text-guided I2I methods [31, 36, 45] do not possess such diversity property.
In Fig. 8, we visually compare our method with related text-guided I2I methods including Null-text inversion (NTI) [31], Prompt-tuning inversion (PTI) [5], PAP [45], FBSDiff [10], and SDEdit [30], as well as ControlNet [48], the representative controllable T2I method. We test different denoising strengths for SDEdit, and test different conditioning modes (Canny edge and depth map) and control weights for ControlNet for comprehensive evaluation of their performance in producing illusion effects. For the remaining methods, we present results with the best illusion effect after extensive hyperparameter tuning. NTI produces results with weak target semantics due to the overtight I2I correlation caused by the attention map consistency constraint, which indicates its deficiency in I2I translation with large semantic deviation. Results of PAP are more aligned to the target semantics but still fall short in text fidelity. Besides, it also suffers from the structural overbinding issue due to direct reusage of reference features during sampling process. Results of PTI and FBSDiff manifest a certain degree of structural-semantic blending, but still underperform in contextual naturalness reflected by visually unpleasant artifacts. Moreover, they suffer from issue of less prominent hidden content, and fail to suit binary text reference images. Results of SDEdit show its difficulty in balancing structural-semantic trade-off, that is, large denoising strength produces results with overwhelmed structural cues while small one is insufficient for the precise translation to the target semantics. Likewise, ControlNet suffers from the same challenge to balance source structure and target semantics when tuning the control weight. By contrast, our PTDiffusion is the only one among the compared methods that realizes harmonious structural-semantic blending, producing visually appealing illusion pictures manifesting both precise textual semantics and clearly discernible hidden content.
@@
Our method adopts DDIM inversion to construct the guidance trajectory (the reconstruction trajectory). To evaluate the necessity of applying DDIM inversion, we have experimented with using the forward diffusion process (i.e., adding noise to $z_{0}$ according to $\hat{z}_{t}\sim\mathcal{N}(\hat{z}_{t};\sqrt{\bar{\alpha}_{t}}z_{0},(1-\bar{\alpha}_{t})\mathcal{I})$) to build the guidance trajectory $\hat{z}_{t}$}, which we term the model w/o inversion. Qualitative results displayed in Fig. 10 show that under the same default hyper-parameters, the generated illusion pictures of our model w/ inversion are superior to the corresponding results produced by the model w/o inversion in both visual quality and structural-semantic blending naturalness. This could be due to that the forward diffusion process introduces randomness (Gaussian noise) to the guidance features, yielding unstable guidance trajectory with feature phase irregularly perturbed at each time step, while the guidance trajectory built upon the DDIM inversion is totally deterministic, providing much more stable phase spectra to be transferred along the sampling process.
Fig. 11 qualitatively demonstrates the effectiveness of our proposed asynchronous phase transfer in controlling the degree of hidden image discernibility. Increasing async distance $d$ enhances visual prominence of hidden content by transferring phase from features at later denoising steps of the reconstruction trajectory into features at earlier denoising steps of the sampling trajectory, while reducing $d$ weakens structural penetration strength by transferring phase from earlier guidance features into later samping features.

### v3

@@ Abstract
Optical illusion hidden picture is an interesting visual perceptual phenomenon where an image is cleverly integrated into another picture in a way that is not immediately obvious to the viewer. Established on the off-the-shelf text-to-image (T2I) diffusion model, we propose a novel training-free text-guided image-to-image (I2I) translation framework dubbed as Phase-Transferred Diffusion Model (PTDiffusion) for hidden art syntheses. PTDiffusion harmoniously embeds an input reference image into arbitrary scenes described by the text prompts, producing illusion images exhibiting hidden visual cues of the reference image. At the heart of our method is a plug-and-play phase transfer mechanism that dynamically and progressively transplants diffusion features’ phase spectrum from the denoising process to reconstruct the reference image into the one to sample the generated illusion image, realizing deep fusion of the reference structural information and the textual semantic information in the diffusion model latent space. Furthermore, we propose asynchronous phase transfer to enable flexible control to the degree of hidden content discernability. Our method bypasses any model training and fine-tuning process, all while substantially outperforming related text-guided I2I methods in image generation quality, text fidelity, visual discernibility, and contextual naturalness for illusion picture synthesis, as demonstrated by extensive qualitative and quantitative experiments. Our project is publically available at this web page.
@@ 1 Introduction
Since diffusion models [21] revolutionizing the field of generative AI, tremendous attention has been focused on various diffusion-based AIGC applications, among which there are also explorations in illusion picture synthesis. For example, DiffQRCoder [29] and Text2QR [47] leverage ControlNet [48] to integrate scannable QR codes into aesthetic pictures. Diffusion Illusions [2] employs T2I diffusion model and score distillation sampling [38, 46] to synthesize images with overlay illusions. Visual Anagrams [16] merges noises estimated from different views to generate multi-view optical illusions, realizing image appearance change under a certain pixel permutation such as image flip, image rotation, or jigsaw rearrangement.
@@
In this paper, we pioneer generating optical illusion hidden pictures (we will use “illusion pictures” as an abbreviation in the following) from the perspective of text-guided I2I translation, i.e., translating an input reference image into an illusion picture that complies with the text prompt in semantic content while manifesting structural visual cues of the reference image. Our goal differs from the aforementioned optical illusion methods in three aspects: (i) different from camouflage image generation [3, 49, 28] that overemphasizes content concealment, we pursue visual discernibility of both target semantic content and hidden visual cues; (ii) unlike synthesizing camouflage image that conceals content into the texture of an existing background image, we expect generating background elements as per the text description; (iii) we do not aim at producing transformation-based (flip, rotation, etc.) optical illusions like Diffusion Illusions [2] and Visual Anagrams [16], but rather focus on seamlessly dissolving a reference image into arbitrary scenes. By contrast, our goal is more methodologically relevant to text-guided I2I [31, 45, 36] and controllable T2I [48, 32, 50] methods. However, since these methods over-bind I2I correlation by explicitly enforcing feature consistency [31, 45, 36] or directly training a control network [48, 32], they are less suitable for I2I translation with large semantic deviation (i.e., suffer from structure-semantic conflict issue), and thus tend to generate contextually unnatural results when applied to synthesize illusion pictures. This enlightens us to explore a disentangled image structure representation to relax I2I correlation binding, as well as an appropriate manner to deeply fuse image structure and semantic information along the sampling process.
@@
Drawing inspiration from digital signal processing that the phase spectrum of an image determines its structural composition, we propose to leverage diffusion features’ phase to disentangle image structure and accordingly propose PTDiffusion, a concise and elegant method based on T2I diffusion model that realizes smooth blending of the reference image’s structural cues and the text-indicated semantic content in the Latent Diffusion Model (LDM) [40] feature space, producing visually appealing illusion pictures in a plug-and-play manner. Specifically, we employ DDIM inversion [43] to construct guidance features along a reference image reconstruction trajectory, and progressively transplant the phase of the guidance features into the corresponding features along the text-guided sampling trajectory, such that structural cues of the input reference image are smoothly penetrated into the sampling process of the target image, yielding generation results exhibiting harmonious illusion effects. Besides, we further propose asynchronous phase transfer to flexibly control the structural penetration strength, endowing our method with controllability to the hidden content discernability in the generated illusion picture. Our method is free from model training, fine-tuning, and online optimization, all while demonstrating noticeable strengths in illusion picture synthesis. The contributions of this work are summarized as follows:
@@ 2 Related work
Text-guided image generation. Since the advent of DDPM [21], diffusion model has soon surpassed GAN [17] on image synthesis [4] and has subsequently been accelerated by DDIM [43] and extended to conditional image generation paradigm by Palette [41]. After large-scale T2I diffusion models [33, 42, 39] remarkably boosting AIGC industry, LDM [40] contributes a classical T2I framework with dramatically lowered computational overhead by transferring DDPM from high-dimensional pixel space into low-dimensional feature space, which inspires subsequent T2I models [37, 7] scaling up to larger capacity. To add controllability to T2I synthesis, ControlNet [48] and T2I-Adapter [32] train a control network of the LDM conditioned on certain image priors (edges, depth maps, etc.), implicitly realizing reference-image-based structural control to the generated images. However, the over-constraining on object contours and shapes of these methods limit their applicability to synthesizing illusion pictures which emphasize harmonious blending of source image structure and target semantics.
Image-to-image translation. I2I translation is initially handled by GAN [17] in a supervised manner [22], and is subsequently extended to unsupervised paradigm through cycle-consistency loss [51], geometry-consistency constraint [9], contrastive learning [35], etc., dramatically enriching image stylization applications such as painting stylization [12, 27, 11], font stylization [1, 23], and image cartoonization [13, 24]. Since the stunning success of T2I synthesis [33, 42, 39, 40], attention has been focused on extending large-scale T2I diffusion models to the realm of text-guided I2I translation. SDEdit [30] translates a reference image by noising it to an intermediate step followed by text-guided denoising. Attention-modulation-based methods such as Null-text inversion [31] and pix2pix-zero [36] correlate source and output image by enforcing consistency of cross-attention maps [18]. Textual-inversion-based methods like Imagic [26] and Prompt Tuning Inversion [5] preserve source image visual information via learnable text embedding. Optimization-free methods represented by PAP [45] maintain source image structure through dynamic feature injection during the reverse sampling process. Besides, FCDiffusion [14] and FBSDiff [10] leverage DCT frequency bands of diffusion features to control different I2I correlation factors. For illusion picture synthesis, however, these methods struggle to produce contextually natural results with both faithful textual semantics and discernable hidden structure due to the overly bound I2I correlation.
@@ 3.1 Overall architecture
To dissolve $x$ into $\tilde{x}$, we propose phase transfer module (PTM) which dynamically blends the structural information of $\hat{z}_{t}$ into $\tilde{z}_{t}$ along the two parallel denoising trajectories. We apply per-step structural penetration realized by PTM only in the early part of the sampling trajectory (which we call phase transfer stage) while leaving the rear part (which we call refining stage) totally unconstrained to guarantee high-quality image synthesis. The two stages are separated by the time step $\lambda T$, where $\lambda$ denotes the proportion of the refining stage to the entire sampling trajectory. The final sampling result $\tilde{z}_{0}$ is transformed to the generated illusion picture via the LDM decoder $D$, i.e., $\tilde{x}=D(\tilde{z}_{0})$.
@@ 3.2 Phase transfer module
where $i$ denotes the imaginary unit, i.e., $i^{2}=-1$. $\hat{R}_{t}$ and $\tilde{R}_{t}$ are the real part of the 2D FFT spectrum of $\hat{z}_{t}$ and $\tilde{z}_{t}$ respectively, $\hat{I}_{t}$ and $\tilde{I}_{t}$ are the corresponding imaginary part. $\hat{M}_{t}$ and $\tilde{M}_{t}$ are the magnitude spectrum of $\hat{z}_{t}$ and $\tilde{z}_{t}$ respectively, $\hat{P}_{t}$ and $\tilde{P}_{t}$ are the corresponding phase spectrum. Then, the extracted phase spectrum $\hat{P}_{t}$ and $\tilde{P}_{t}$ are linearly blended with a time-dependent blending coefficient $b_{t}$, yielding the fused phase spectrum $P_{t}^{fuse}$ as follows:
@@
The fused phase is recombined with the original magnitude $\tilde{M}_{t}$ before transforming back to the spatial domain with 2D IFFT, finally resulting in the structurally penetrated $\tilde{z}_{t}$:
@@
Since the structural information of the guidance features $\hat{z}_{t}$} is becoming increasingly prominent as the denoising process proceeds, direct phase transfer in the later denoising steps is prone to harm contextual naturalness of the final result due to excessive structural penetration. Thus, we design a decayed phase blending schedule which gradually decays the blending coefficient {$b_{t}$} in the later denoising steps of the phase transfer stage, which we formulate as below:
@@
where $\lambda\leq\tau\leq 1$. It means that we apply direct phase replacement in the early part of the phase transfer stage where the guidance features $\hat{z}_{t}$ are structurally less distinct, while gradually decaying phase transfer intensity in the later section of the phase transfer stage as $\hat{z}_{t}$ becomes structurally more and more prominent. By decaying phase transfer intensity to avoid overly penetrated structural information, both the visual quality and the contextual naturalness of the generated illusion picture can be noticeably improved.
@@ 3.3 Asynchronous phase transfer
Since sufficient denoising steps in the refining stage is of crucial importance to ensure image quality, we fix $\lambda$, namely the length of the phase transfer stage, and propose asynchronous phase transfer to realize controllable structural penetration strength within the fixed-length phase transfer stage. As Fig. 4 displays, later $\hat{z}_{t}$ in the reconstruction trajectory is structurally more prominent than its earlier counterpart. This inspires us to transfer phase from later $\hat{z}_{t}$ with smaller time steps into earlier $\tilde{z}_{t}$ with larger time steps to enhance structural penetration, namely the so-called asynchronous phase transfer as illustrated in Fig. 5.
@@
To this end, we design a concise and elegant solution that implements asynchronous phase transfer based on simple synchronous denoising, i.e., parallel denoising along the reconstruction and sampling trajectory. Specifically, given an async distance $d$, we firstly leverage $\hat{z}_{t}$ to estimate its future counterpart $\hat{z}_{t-d}$ that is $d$ steps ahead of itself in the reconstruction trajectory, then transfer the phase spectrum of the estimated $\hat{z}_{t-d}$ into $\tilde{z}_{t}$. Let APTM denote the asynchronous phase transfer module, we formulate it as follows:
@@ 4.1 Qualitative results
As demonstrated in Fig. 7, despite using deterministic DDIM inversion and sampling, our method allows to sample diverse results simply by varying the initialized Gaussian noise $\tilde{z}_{T}$, while existing advanced text-guided I2I methods [31, 36, 45] do not possess such diversity property.
In Fig. 8, we visually compare our method with related text-guided I2I methods including Null-text inversion (NTI) [31], Prompt-tuning inversion (PTI) [5], PAP [45], FBSDiff [10], and SDEdit [30], as well as controllable T2I method represented by ControlNet [48]. We test different denoising strengths for SDEdit, and test different conditioning modes (Canny edge and depth map) and control weights for ControlNet for comprehensive evaluation of their performance in producing illusion effects. For the remaining methods, we present results with best illusion effect after hyperparameter tuning. NTI produces results with weak target semantics due to the overtight I2I correlation caused by the attention map consistency constraint, which indicates its deficiency in I2I translation with large semantic deviation. Results of PAP are more aligned to the target semantics but still fall short in text fidelity. Besides, it also suffers from the structural overbinding issue due to direct reusage of reference features during the sampling process. Results of PTI and FBSDiff manifest a certain degree of structure-semantic blending, but still underperform in contextual naturalness reflected by visually unpleasant artifacts. Moreover, they suffer from issue of less prominent hidden content, and fail to suit binary text reference images. Results of SDEdit show its difficulty in balancing structure-semantic trade-off, i.e., large denoising strength produces results with overwhelmed structural cues while small one is insufficient to translation input image to the target semantics. Likewise, ControlNet suffers from the same challenge to balance source structure and target semantics when tuning the control weight. By contrast, our PTDiffusion is the only one among the compared methods that realizes harmonious structure-semantic blending, producing visually appealing illusion pictures manifesting both precise textual semantics and clearly discernible hidden content.
@@
Our method adopts DDIM inversion to construct the guiding trajectory (the reconstruction trajectory). To verify the necessity of applying DDIM inversion, we have experimented with using the forward diffusion process (i.e., adding noise to $z_{0}$ according to $\hat{z}_{t}\sim\mathcal{N}(\hat{z}_{t};\sqrt{\bar{\alpha}_{t}}z_{0},(1-\bar{\alpha}_{t})\mathcal{I})$) to build the guiding trajectory $\hat{z}_{t}$}, which we term the model w/o inversion. Qualitative results (evaluated under the same default hyper-parameters) displayed in Fig. 10 show that the illusion pictures generated by our model w/ inversion are superior to the corresponding results w/o inversion in both visual quality and structure-semantic blending naturalness. This could be due to that the forward diffusion process introduces randomness (Gaussian noise) to the guidance features, yielding unstable guiding trajectory with feature phase irregularly perturbed at each time step, while the guiding trajectory built with DDIM inversion is totally deterministic, providing more stable phase spectra to be transferred along the sampling process.
Fig. 11 qualitatively demonstrates the effectiveness of our proposed asynchronous phase transfer in controlling hidden image discernibility. Increasing async distance $d$ enhances hidden content visual prominence by transferring phase from features at later denoising steps of the reconstruction trajectory into features at earlier denoising steps of the sampling trajectory, while reducing $d$ weakens hidden content discernibility by conversely transferring phase from earlier guidance features into later samping features.